OpenStudy (anonymous):

Solve the following equation for all solutions of x. 4x(squared) - 49 = 0 A. x=-2/7, 2/7 B. x= 0, 7/2 C. x= -7/2, 7/2 D. x= -7/2

OpenStudy (hoblos):

4x^2 - 49=0 (2x)^2 - 7^2 =0 (2x-7)(2x+7)=0 x= 7/2 , x= -7/2

OpenStudy (hoblos):

C

OpenStudy (anonymous):

But I don't understand how you got the answer from (2x - 7)(2x + 7) = 0.

OpenStudy (hoblos):

2x+7=0 OR 2x-7=0 => x=-7/2 => x=7/2

OpenStudy (anonymous):

Oh, I see.....

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

my pleasure XD